Simplify to a single trig function:
(cos3x-cosx)/
(sin3x-sinx)
PLEASE HELP ME I DON'T KNOW WHAT TO DO!!!

Answer by stanbon(48558)

Working with just the cos(3x) you get the following:
cos3x = cos(x)cos(2x) - sin(x)sin(2x)
=cos(x)[1-2sin^2(x)] - sin(x)[2sin(x)cos(x)]
=cos(x)-2cos(x)sin^2(x) - 2sin^2(x)cos(x)
=cos(x) - 4sin^2(x)cos(x)
-----------------------------
So subtracting cos(x) from this result you would get:
= - 4sin^2(x)cos(x)
==========================
Cheers,
Stan H.